# A Smooth Approximation Framework for Weakly Convex Optimization

### v1

@@ Appendix B Missing Proofs
$\lVert\mathbf{x}^{k}-\hat{\mathbf{x}}^{k}\rVert^{2}\leq\lVert\mathbf{x}^{k-1}-\hat{\mathbf{x}}^{k}\rVert^{2}+\frac{1}{\hat{\rho}-\bar{\rho}+\gamma^{-1}}\left[\frac{\lVert\bm{\delta}^{k-1}\rVert^{2}}{(\gamma^{-1}-\hat{\rho}-L_{\eta})}-2\langle\bm{\delta}^{k-1},\mathbf{x}^{k-1}-\hat{\mathbf{x}}^{k}\rangle-2(\hat{\rho}-\bar{\rho})\lVert\mathbf{x}^{k-1}-\hat{\mathbf{x}}^{k}\rVert^{2}\right].$
@@
$\phi_{\eta}^{\hat{\rho}}(\mathbf{x}^{k})$
$=\min_{\mathbf{x}}\,\big\{\phi_{\eta}(\mathbf{x})+\frac{\hat{\rho}}{2}\lVert\mathbf{x}-\mathbf{x}^{k}\rVert^{2}\big\}$
@@
$\leq\phi_{\eta}(\hat{\mathbf{x}}^{k})+\frac{\hat{\rho}}{2}\lVert\hat{\mathbf{x}}^{k}-\mathbf{x}^{k-1}\rVert^{2}+\frac{\hat{\rho}}{2(\hat{\rho}-\bar{\rho}+\gamma^{-1})}\left[\frac{\lVert\bm{\delta}^{k-1}\rVert^{2}}{(\gamma^{-1}-\hat{\rho}-L_{\eta})}-2\langle\bm{\delta}^{k-1},\mathbf{x}^{k-1}-\hat{\mathbf{x}}^{k}\rangle-2(\hat{\rho}-\bar{\rho})\lVert\hat{\mathbf{x}}^{k}-\mathbf{x}^{k-1}\rVert^{2}\right]$
@@
Suppose we take $\gamma=(c\sqrt{K}+\hat{\rho}+L_{\eta})^{-1}$ where $c=\sqrt{\frac{\hat{\rho}}{(\Delta+R\eta)m}}\sigma$, then we have
@@
$=\frac{\hat{\rho}}{\hat{\rho}-\bar{\rho}}\left\{\frac{(2\hat{\rho}-\bar{\rho}+L_{\eta})(\Delta+R\eta)}{K}+\sqrt{\frac{\hat{\rho}(\Delta+R\eta)}{mK}}\sigma\right\}$
@@
$\langle\nabla g(\mathbf{x}^{\star}),\mathbf{x}^{\star}-\mathbf{z}^{t}\rangle+\pi(\mathbf{x}^{\star}\rangle-\pi(\mathbf{z}^{t})\leq 0.$
@@
$\leq\frac{\gamma}{2}\lVert\mathbf{x}^{\star}-\mathbf{z}^{t-1}\rVert^{2}+\bar{L}D\lVert\mathbf{z}^{t}-\mathbf{x}^{\star}\rVert,$ (86)
where the last inequality follows from the Lipschitz continuity of $\nabla g$ and the assumption $\lVert\mathbf{x}^{\star}-\mathbf{x}^{t}\rVert\leq D$. Define $D_{s}=\lVert\mathbf{x}^{\star}-\mathbf{z}^{s}\rVert$, and we denote $\bar{L}=\sup\{L(\mathbf{x},\mathbf{y}):\mathbf{x},\mathbf{y}\in\mathcal{B}_{D}(\mathbf{x}^{\star})\}$. Then, applying (86), we have
@@
$\lVert\mathbf{y}^{t}-\mathbf{x}^{\star}\rVert$ $=(1-\alpha_{t})\lVert\mathbf{y}^{t-1}-\mathbf{x}^{\star}\rVert+\alpha_{t}\lVert\mathbf{z}^{t}-\mathbf{x}^{\star}\rVert$
@@
$\leq\left(1+\frac{2\bar{L}}{\hat{L}}\right)D,$ (87)
@@
We claim that the line search will terminate in at most $k^{\star}=\max\left\{\log_{\tau_{\mathrm{u}}}\frac{2\tilde{L}}{\bar{L}_{t}},0\right\}+1$ steps. To see this, suppose for the sake of contradiction that the line search does not terminate within $k^{\star}$ iterations. Then, upon reaching $\hat{L}=\bar{L}_{t}\tau_{\mathrm{u}}^{k^{\star}+1}$, we would have $\hat{L}\geq 2\tilde{L}$. By applying inequality (87), we observe that

### v3

@@ Appendix B Missing Proofs
$\lVert\mathbf{x}^{k}-\hat{\mathbf{x}}^{k}\rVert^{2}\leq\lVert\mathbf{x}^{k-1}-\hat{\mathbf{x}}^{k}\rVert^{2}+\frac{1}{\hat{\rho}-\bar{\rho}+\gamma^{-1}}\Big[\frac{\lVert\bm{\delta}^{k-1}\rVert^{2}}{(\gamma^{-1}-\hat{\rho}-L_{\eta})}-2\langle\bm{\delta}^{k-1},\mathbf{x}^{k-1}-\hat{\mathbf{x}}^{k}\rangle-2(\hat{\rho}-\bar{\rho})\lVert\mathbf{x}^{k-1}-\hat{\mathbf{x}}^{k}\rVert^{2}\Big].$
@@
$\phi_{\eta}^{\hat{\rho}}(\mathbf{x}^{k})$ $=\min_{\mathbf{x}}\,\big\{\phi_{\eta}(\mathbf{x})+\frac{\hat{\rho}}{2}\lVert\mathbf{x}-\mathbf{x}^{k}\rVert^{2}\big\}$
@@
$\leq\phi_{\eta}(\hat{\mathbf{x}}^{k})+\frac{\hat{\rho}}{2}\lVert\hat{\mathbf{x}}^{k}-\mathbf{x}^{k-1}\rVert^{2}+\frac{\hat{\rho}}{2(\hat{\rho}-\bar{\rho}+\gamma^{-1})}\Big[\frac{\lVert\bm{\delta}^{k-1}\rVert^{2}}{(\gamma^{-1}-\hat{\rho}-L_{\eta})}$-
$\qquad 2\langle\bm{\delta}^{k-1},\mathbf{x}^{k-1}-\hat{\mathbf{x}}^{k}\rangle-2(\hat{\rho}-\bar{\rho})\lVert\hat{\mathbf{x}}^{k}-\mathbf{x}^{k-1}\rVert^{2}\Big]$
@@
Suppose we take $\gamma=(c\sqrt{K}+\hat{\rho}+L_{\eta})^{-1}$ where $c=\sqrt{\frac{\hat{\rho}}{2(\Delta+R\eta)m}}\sigma$, then we have
@@
$=\frac{\hat{\rho}}{\hat{\rho}-\bar{\rho}}\left\{\frac{(2\hat{\rho}-\bar{\rho}+L_{\eta})(\Delta+R\eta)}{K}+\sqrt{\frac{2\hat{\rho}(\Delta+R\eta)}{mK}}\sigma\right\},$
@@
$\langle\nabla g(\mathbf{x}^{\star}),\mathbf{x}^{\star}-\mathbf{z}^{t}\rangle+\pi(\mathbf{x}^{\star})-\pi(\mathbf{z}^{t})\leq 0.$
@@
$\leq\frac{\gamma}{2}\lVert\mathbf{x}^{\star}-\mathbf{z}^{t-1}\rVert^{2}+\bar{L}D\lVert\mathbf{z}^{t}-\mathbf{x}^{\star}\rVert,$
where the last inequality follows from the Lipschitz continuity of $\nabla g$ and the assumption $\lVert\mathbf{x}^{\star}-\mathbf{x}^{t}\rVert\leq D$. Define $D_{s}=\lVert\mathbf{x}^{\star}-\mathbf{z}^{s}\rVert$, and we denote $\bar{L}=\sup\{L(\mathbf{x},\mathbf{y}):\mathbf{x},\mathbf{y}\in\mathcal{B}_{D}(\mathbf{x}^{\star})\}$. Then, applying (B.8), we have
@@
$\lVert\mathbf{y}^{t}-\mathbf{x}^{\star}\rVert$ $\leq(1-\alpha_{t})\lVert\mathbf{y}^{t-1}-\mathbf{x}^{\star}\rVert+\alpha_{t}\lVert\mathbf{z}^{t}-\mathbf{x}^{\star}\rVert$
@@
$\leq\left(1+\frac{2\bar{L}}{\hat{L}}\right)D,$
@@
We claim that the line search will terminate in at most $k^{\star}=\max\left\{\log_{\tau_{\mathrm{u}}}\frac{2\tilde{L}}{\bar{L}_{t}},0\right\}+1$ steps. To see this, suppose for the sake of contradiction that the line search does not terminate within $k^{\star}$ iterations. Then, upon reaching $\hat{L}=\bar{L}_{t}\tau_{\mathrm{u}}^{k^{\star}+1}$, we would have $\hat{L}\geq 2\tilde{L}$. By applying inequality (B.8), we observe that